If A is of 3 rd order square matrix such that |A|=2 then |adj(adj(adjA))| is=? Pls ans with proper steps

If A is of 3rd order square matrix such that |A|=2 then |adj(adj(adjA))| is=?
Pls ans with proper steps

Grade:12

1 Answers

Arun
5 years ago
Dear Manan
 
As we know
|adj A| = |A|n –1
Now
|adj(adj(adjA))| = |(adj(adjA))|n –1 = |(adjA))|(n –1)^2 = |A|(n –1)^3
 = 22^3
 = 28
 = 256
 
Regards
Arun (askIITians forum expert)